Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2024 logbook, Contact: F.Wolff-Fabris, A.Galler, H.Sinn  Not logged in ELOG logo
Message ID: 148     Entry time: 12 Aug 2024, 23:47
Author: Raphael de Wijn 
Group: SPB 
Subject: Shift summary 

X-ray delivery

11.56 keV, 202 pulses @0.5 MHz, ~0.9 mJ, stable delivery

Optical laser delivery

800 ns laser for jet visualization
PPLaser for data collection

Achievements/Observations

  • Finished in-house PYP data collection
  • Tested in-house GFP sample
  • 2 hotfixes for the online calibration pipeline in the morning

Issues

Last PCS beamtime cancelled as the sample did not survive the shipping from US

ELOG V3.1.4-7c3fd00